OpenStudy (moonlitfate):

Help with derivatives; everyone loves those, right? Find the derivative of the function: f(x) = (x^3 - 6x^2 + 5)/x^2

OpenStudy (moonlitfate):

So far this is what I've done: rewritten the equation to look like this: x^3 - 6x^2 +5x^-2

OpenStudy (raden):

use the division rule : y = u/v y' = (u'v - uv')/v^2

OpenStudy (moonlitfate):

And tried to find the derivative, but i feel like I'm missing something; I have this so far: f'(x) = (3x^2 - 12x -10)/x^3

OpenStudy (raden):

in this case, u = x^3 - 6x^2 + 5 ----> u'= ... ? v = x^2 ----> v' = ... ?

hartnn (hartnn):

in just an alternate way, we can avoid quotient rule...

OpenStudy (moonlitfate):

Oh... just do them piece by piece...

OpenStudy (raden):

or u can divided all things in numerator by x^2 so, (x^3 - 6x^2 + 5)/x^2 = x^3/x^2 - 6x^2/x^2 + 5/x^2 simplify again

OpenStudy (moonlitfate):

u' = 3x^2 -12x v' = 2x

OpenStudy (raden):

yes, that's right if u want use the division rule apply it to formula above

OpenStudy (moonlitfate):

@RadEn-- After applying that formula, I got the answer: f'(x) = (3/2)x -6

OpenStudy (raden):

let's check it ... y' = (u'v - uv')/v^2 y' = [(3x^2 -12x)x^2 - (x^3-6x^2+5)(2x)]/(x^2)^2 y' = (3x^4 - 12x^3 - 2x^4 + 12x^3 - 10x)/x^4 y' = (x^4 - 10x)/x^4 y' = x^4/x^4 - 10x/x^4 y' = 1 - 10/x^3 y' = 1 - 10x^(-3) hope i didnt mistake

OpenStudy (raden):

wolfram has the answer like m ;) http://www.wolframalpha.com/input/?i=f%28x%29+%3D+%28x^3+-+6x^2+%2B+5%29%2Fx^2